I have an unknown volume of gas at a pressure of 0.50 atm and a temperature of 325 K. If I raise the pressure to 1.2 atm, decrease the temperature to 320 K, and measure the final volume to be 48 liters, what was the initial volume of the gas? (round to correct sig figs)


1
Expert's answer
2021-08-23T02:56:48-0400

Assuming that the number of mols are constant for both conditions:

(P1 x V1)/T1 = (P2 x V2)/T2

Now you plug in the given values. V_1 is the unknown.

(0.5 x V1)/325 = (1.2 x 48)/230

Separate V_1

V1 = (1.2 x 48 x 325)/ (230 x 0.5) = 162.782608696 L 

There are 2 sig figs

V= 160 L